Write the equation in logarithmic form, show/explain work.
64=2^6

Found 2 solutions by sachi, RedemptiveMath:
Answer by sachi(548)

64=2^6
here 2 is the base & 6 is exponent
so in log form log 64=6
2
i.e. log of 64 to base 2 is 6
ans

Answer by RedemptiveMath(80)

The standard form for an exponential function is given as x = b^y. In our example, 64 corresponds to x, 2 corresponds to b, and 6 corresponds to y. The explanations of the letters are given the in logarithm explanation.

Standard logarithmic form is given as logb (x) = y, where b is the base (written in subscript form), x is the "product of factors" or answer to an exponential function, and y is the exponent used in the exponential function. A logarithm is basically the inverse of a exponential function. Now that we have this information, we can convert the exponential function into a logarithmic function:

64=2^6
log2 (64) = 6.

Make sure that 2 is written as a subscript. The logarithmic function uses the base and the product of factors to find the exponent, while the exponential function uses the base and exponent to find the product of factors. We can try to find logarithms by asking, "The base to what power equals the product of factors?" Or, "2 to what power is 64." The answer is 6.